OpenStudy (mony01):

Decide whether the set S is a linearly independent subset of the given vector space V. V=R^3 S=([2,-1,4],[4,-2,8])

OpenStudy (anonymous):

[2,-1,4] = (1/2) (4,-2,8] . So no. S only span a line in R^3

OpenStudy (mony01):

how did you get 1/2?

OpenStudy (mony01):

so its dependent?

OpenStudy (kirbykirby):

Yes the vectors are dependent since one of them is a scalar multiple of the other one

OpenStudy (kirbykirby):

You can view it in either way: [2,-1,4] = (1/2)* [4,-2,8] or (2)*[2,-1,4] = [4,-2,8]